Laylaheck:

Jill has $8.75. Jack has three times as much as Jill, but he spent $5.00 to buy a book. How much money does Jack have?

Vocaloid:

this is my interpretation let x = Jack and y = Jill x - 5 is Jack's money after he paid for the book x - 5 = 3 * y solve for x, given that y = 8.75